Find two numbers with a sum of –32 and a difference of 2
–18 and –16
17 and 15
–16 and –14
–17 and –15

Answer by Edwin McCravy(20055)

Find two numbers with a sum of –32 and a difference of 2

-----------------------------------------------

–18 and –16   

       sum = (-18)+(-16) = -34
difference = (-18)-(-16) = (-18)+(16) = -2

-----------------------------------------------

 17 and  15

       sum = (17)+(15) = 32

difference = (17)-(15) = (17)+(-15) = 2

-----------------------------------------------

–16 and –14

       sum = (-16)+(-14) = -30
difference = (-16)-(-14) = (-16)+(14) = -2

-----------------------------------------------

–17 and –15

       sum = (-17)+(-15) = -32
difference = (-17)-(-15) = (-17)+(15) = -2

-----------------------------------------------
Sorry, none of the choices have both a sum of -32 and 
also a difference of 2.

If the last one had been given in the opposite order, 
it would have been the answer:  


–15 and –17

       sum = (-15)+(-17) = -32
difference = (-15)-(-17) = (-15)+(17) = 2

But they were not given in that order.  Any time a problem states
"the difference of two numbers, it always means the FIRST number 
listed MINUS the SECOND number listed.  You can't reverse them.

Edwin